OpenStudy (anonymous):

A storeowner buys a case of 144 pens for $28.80. Tax and shipping cost an additional $8.64. He sells the pens for $.59 each. What is the markup per pen? What is the percent of markup?

OpenStudy (mertsj):

How much does he have invested in each pen including the tax and shipping?

OpenStudy (anonymous):

28.80 + 8.64= 37.44

OpenStudy (mertsj):

That is for all 144 pens. How much for each pen?

OpenStudy (anonymous):

He is selling each for .59 each

OpenStudy (mertsj):

How much did he pay for each pen?

OpenStudy (anonymous):

Im not sure what to do

OpenStudy (mertsj):

Divide the total cost by the number of pens.

OpenStudy (anonymous):

.26

OpenStudy (anonymous):

so he paid .26 for each pen which is a .33 markup

OpenStudy (mertsj):

And the percent markup is the markup over the original price multiplied by 100

OpenStudy (anonymous):

so 55.9% or round to 56% ?

OpenStudy (mertsj):

.33/.26=33/26=1.27=127% markup You know he more than doubled the price so the markup has to be more than 100 % of what he paid.

OpenStudy (anonymous):

OK I see thank you